OpenStudy (dannyrod2000):

2log2 15-log2 5+log2 3

OpenStudy (dannyrod2000):

\[2\log _{2}15-\log _{2}5+\log _{2}3\]

OpenStudy (dannyrod2000):

which of the following is equivalent to A) 2log2 9 B) 2log2 10 C) log2 18 D) log2 135

OpenStudy (dannyrod2000):

@JoannaBlackwelder

OpenStudy (dannyrod2000):

ur fast @JoannaBlackwelder

OpenStudy (joannablackwelder):

Lol, just saw it. :-) Do you have any thoughts on how to start?

OpenStudy (dannyrod2000):

idk i put it on wolfram alpha

OpenStudy (joannablackwelder):

Ok, so did that show you the related roperties?

OpenStudy (joannablackwelder):

*properties

OpenStudy (dannyrod2000):

yes it did

OpenStudy (dannyrod2000):

alternate forms

OpenStudy (dannyrod2000):

it gave me \[\log(135)/\log(2)\]

OpenStudy (joannablackwelder):

Yeah, that is true. Can you tell which answer that is?

OpenStudy (dannyrod2000):

ohh i know what to do.

OpenStudy (dannyrod2000):

give me a sec

OpenStudy (sklopr):

It says you are a human calculator

OpenStudy (dannyrod2000):

they both are equal to 7.07

OpenStudy (dannyrod2000):

am i right??

OpenStudy (welshfella):

= log2 15^2 - log2 5 + log2 3 = log2 225/5 + log2 3 = log2 45 + log2 3 - can you finish this?

OpenStudy (joannablackwelder):

I agree with D. :-)

OpenStudy (dannyrod2000):

thanks for the help!

OpenStudy (sklopr):

If you are a Human Calculator you should be able to do this

OpenStudy (joannablackwelder):

You're welcome. :-)

OpenStudy (dannyrod2000):

yes @welshfella its 7.07

OpenStudy (welshfella):

thers no need to write it as a number the answer is D log2 135

OpenStudy (dannyrod2000):

yes i know i got that

OpenStudy (dannyrod2000):

its easier for me in number form